Hi
the decimal ratio of the number of share price rises to falls is 2.7.
if 5000 share prices of companies have fallen, find
the number of share prices which have risen
the total number of listed companies.
thanks

Answer by Theo(13342)

r is the number of rises.
f is the number of falls.

if the ratio of rises to falls is 2.7, then you get:

r/f = 2.7/1

solve for r to get r = 2.7 * f

if f = 5000, then r = 2.7 * 5000 = 13500.

the number of share prices that have risen is 13500.

13500 / 5000 = 2.7

the total number of listed companies would be at least 13500 + 5000 = 18500.

this does not include any companies whose share prices have remained the same.

the assumption is that all the listed companies have either had their share prices rising or falling or remaining the same.